PLEASE HELP!! I just need the system of equations and it says there are only four but I think there might be more. My teacher is strict and has a strong accent so I can’t always understand or ask for help.
The International Canine Academy raises and trains Siberian sled dogs and dancing French poodles.Breeders supply at most 20 poodles and 15 sled dogs each year. Each poodle eats 2lb of food a day.Each sled dog eats 6lb of food per day. Supplies restricted the academy to 100lb of food per day. Poodles require 1,000 hr/yr training, whereas sled dogs require 250hr/yr.The academy cannot provide more than 15,000 hr/ur of training.

Answer by Theo(13342)

let x = the number of french poodles
let y = the number of siberian sled dogs

the breeders supply at most 20 poodles and 15 sled dogs each year.
your constraint functions to model that are:

x <= 20
y <= 15

each poodle eats 2 pounds of food each day and each sled dog eats 6 pounds of food each day.
the amount of food available per day is restricted to 100 pounds.
your constraint function to model that is.

2x + 6y <= 100

each poodle requires 1000 hours of training per year and each sled dog requires 250 hours of training per year.
the number of hours available for training each year cannot be more than 15,000 hours.
your constraint function to model that is:

1000x + 250y <= 15000

the number of poodles and the number of sled dogs must be greater than or equal to 0.
your constraint functions to model that are:

x >= 0
y >= 0

putting them all together, your constraint functions are:

x <= 20
y <= 15
2x + 6y <= 100
1000x + 250y <= 15000
x >= 0
y >= 0

those constraint functions are your solution as best i can determine.